There are a total of 100 fishes in a pond for a fishing tournament. 25 are bluegill, 25 are largemouth bass, 35 are channel catfish, and 15 are crappies.

Question A: What is the probability of someone catching bluegill, then a catfish, then bluegill and finally a crappy without replacing the fish back to the pond.

Question B: What is the probability someone catching a crappy back to back if after each catch the fish is put back to the pond.
Answer #1

A)probability of someone catching bluegill, then a catfish, then bluegill and finally a crappy without replacing the fish back to the pond=(25/100)*(35/99)*(24/98)*(15/97)=0.003347

B)

probability someone catching a crappy back to back if after each catch the fish is put back to the pond

=(15/100)*(15/100) =0.0225